The standard enthalpy of vaporization (Deltavap) of methylacrylate (CH2CHCO2CH3) is 36.0 kJ/mol.
The normal boiling point of methylacrylate is 80.5°C.  
Calculate the vapour pressure of methylacrylate at 43.0°C.
Answer #1

A( - In R Pi = 4 atm at T 80-5+}3 k T= 353.5 P2=? at T2=43 +213= 316 K From In -36X10 m 1 316 8314 mo1s In (무) 1.453 2 303 l0